# A flexible class of latent variable models for the analysis of antibody response data

### v2

@@ 1 Introduction
Serological data analysis plays a central role in reconstructing individual- and population-level exposure histories to infectious diseases, using information derived from antibody concentration measurements (metcalf2016; corran2007). Antibody levels can function as biomarkers of prior exposure, enabling the detection of both recent infections and long-term immune responses. In low-prevalence or post-elimination settings, serological data become particularly valuable as they can be used to identify subtle signs of disease resurgence or reintroduction (Drakeley2005; arnold2018). Consequently, serological surveillance offers a powerful approach for monitoring disease dynamics, especially where traditional diagnostic methods may be insufficiently sensitive due to sporadic infections or predominantly asymptomatic transmission (Drakeley2005).
Gaussian mixture models (GMMs) are among the most widely used methods for analysing antibody concentration data (fraley2002; arnold2018). In the serological context, a GMM typically assumes that the antibody concentration $y_{i}$ for the $i$-th individual arises from a mixture of two Gaussian distributions, representing seronegative and seropositive subpopulations:
@@
where $\pi_{0}+\pi_{1}=1$, and $(\mu_{k},\sigma_{k}^{2})$ denote the mean and variance of each component. This binary specification reflects the assumption that individuals fall into one of two mutually exclusive serological states, with no allowance for intermediate or uncertain exposure histories. Across a wide range of pathogens, such mixture models have become a standard way to operationalise immune activation from serological measurements by separating individuals into low- and high-response groups interpreted as immunologically naïve versus previously exposed or vaccinated. Although GMMs provide a flexible, data-driven approach to defining seropositivity thresholds, alternative strategies are also common in sero-epidemiology (Hay2024Serodynamics): using manufacturer-defined assay cutoffs; determining thresholds via ROC curves when positive and negative controls are available; basing cutoffs on the distribution of negative controls alone (typically two or three standard deviations above the mean); or, when available, applying immunological correlates of protection to define seroprotection-based thresholds.
@@
This approach builds on an extensive body of work applying GMMs to diverse pathogens for which serology provides a reliable marker of past exposure, including directly transmitted childhood infections such as measles, mumps and rubella (Vyse2006); parasitic and neglected tropical diseases such as onchocerciasis (Golden2016) and trachoma (Migchelsen2017); mosquito-borne infections such as malaria, where mixture models underpin estimates of age-specific seroprevalence (Sepulveda2015); and viral infections characterised by substantial asymptomatic transmission such as hepatitis E (Katuwal2024) and SARS-CoV-2 (Bottomley2021). These applications highlight that the usefulness of serology strongly depends on the biology of the pathogen. As emphasised by metcalf2016, serological surveys are most informative for infections in which antibody responses provide a durable and interpretable record of prior exposure or immunity. This includes infections that generate long-lasting protection, those with measurable antigenic variation, and those in which antibodies reliably reflect cumulative exposure even when infections are predominantly asymptomatic or clinically under-ascertained. The methods presented in this paper are relevant for the analysis of these types of pathogens.
@@
Recent methodological advances have sought to extend GMMs by incorporating covariates, such as age or other demographic factors, which influence antibody concentrations and seropositivity status. For example, hardelid2008 used maternal age and country-of-birth to model variation in rubella antibody levels among newborns using a finite-mixture regression framework. Age-dependent mixture models are especially common in infectious disease applications: delfava2016 embedded a three-component mixture within a catalytic model to estimate age-specific immunity to varicella-zoster virus, while cox2022 used penalised splines to let seropositivity probabilities vary smoothly with age in dengue serology data. Similarly, kyomuhangi2021 combined antibody acquisition and catalytic models to allow for temporally varying seroconversion and boosting rates. These approaches reflect cumulative exposure risk by allowing mixing proportions or transition probabilities to evolve with age or other covariates. In parallel, efforts have been made to relax the normality assumption within each component. A key extension replaces Gaussian components with skew-normal distributions (azzalini1985), enabling the model to capture asymmetries commonly observed in antibody concentration data. Skew-normal mixtures retain analytical tractability while offering greater flexibility in representing the shape of seronegative and seropositive subpopulations (dias_domingues2024).
@@
An alternative line of work has sought to avoid explicit dichotomisation through antibody acquisition (AA) models, in which age-dependent boosting and waning processes generate smooth trajectories of mean antibody levels (yman2016). Although these models provide a useful mechanistic link between age, exposure, and antibody dynamics, they typically rely on a log-Gaussian assumption and do not adequately accommodate the skewness and multimodality that characterise many serological datasets. Semiparametric methods have also been proposed that use ensemble machine learning to estimate age-antibody curves nonparametrically (arnold2017), but these target only the mean antibody level, precluding inference on other parameters of epidemiological interest. The proposed framework in this study unifies the strengths of these approaches by accommodating both mechanistic interpretability through AA-type formulations and flexible modeling of the full heterogeneous distribution of antibody levels across age through a latent variable formulation.
@@
A key practical challenge in implementing the proposed framework is the computational burden associated with maximum likelihood estimation, which requires repeated numerical integration. To address this, we develop a computationally efficient alternative based on histogram-based $L_{2}$ minimum distance estimation that substantially reduces the computational burden. Our work builds on the $L_{2}$ error theory developed by scott2001parametric, which establishes the criteria under which $L_{2}$ estimators converge to the true parameter values. We extend this framework to our latent variable modelling framework and establish consistency of the resulting estimator under appropriate regularity conditions.
@@
The remainder of the paper is organised as follows. Section 2 introduces the proposed latent variable modelling framework, contrasts it with alternative latent formulations, and presents both single-density and mixture specifications for the latent seroreactivity distribution. Section 3 describes how age-dependent structure can be incorporated through mechanistic and data-driven parameterisations of both single Beta distributions and finite mixtures. Section 4 details maximum likelihood and histogram-based $L_{2}$ estimation procedures, establishes the consistency of the $L_{2}$ estimator, and describes model validation approaches. Section 5 presents a comprehensive simulation study comparing the statistical and computational efficiency of $L_{2}$ and maximum likelihood estimation across epidemiologically relevant scenarios. Section 6 demonstrates the framework through applications to malaria serology data for AMA1 and MSP1 antigens, including age-stratified model comparisons. We conclude in Section 7 with methodological implications and extensions to multivariate settings, spatial-temporal modelling, and high-dimensional omics applications.
@@ 2 The proposed latent variable modelling framework for serological data analysis
To illustrate the interpretation, let us consider the simplest formulation of antibody acquisition models originally proposed by yman2016. According to this model, we define
@@ 3.2 Finite mixture models with age dependent mixing probabilities
A wide range of mechanistic models for the relation between age and antibody levels has been developed in the context of infectious disease research. These include catalytic models, models that allow repeated boosting through multiple exposures, and models with time varying hazards; a comprehensive overview of these approaches can be found in Hay2024Serodynamics. Within our latent mixture framework, these formulations offer a convenient way to specify $\pi(a)$ while retaining epidemiological interpretability. For example, the catalytic framework assumes that the cumulative probability of activation follows
@@ 4 Inference
Parameter estimation for the latent variable model requires integrating over the continuous distribution of the unobserved immune state $T$, which presents computational challenges for likelihood-based inference. We address this by considering and comparing two approaches: standard maximum likelihood estimation (MLE), and a computationally efficient histogram-based $L_{2}$ distance minimization method. This can be thought of a discretizations of the $L_{2}E$ method proposed by scott2001parametric. Although our main motivation for using the method is computational gain, the limiting case i.e the $L_{2}E$ method is also used for it’s robusntess scott2001parametric.
@@
The theoretical justification for this approach rests on the fact that, as the bin widths $\Delta_{j}$ shrink, the histogram density $\widehat{f}_{j}$ converges to the empirical distribution of the data, while $f(m_{j};\vartheta,\psi)$ converges to the model density at those points. In this limiting regime, minimisation of (13) defines a minimum–distance estimator (wolfowitz1957; beran1977) that minimises the $L_{2}$ distance between the empirical and model densities, i.e. the $L_{2}E$ method.
@@ 4.1 Consistency of the histogram-based $L_{2}$ estimator
We establish the consistency of our estimator using the asymptotic framework of $M$-estimation (vandervaartwellner1996). The estimator is defined by minimizing the integrated squared error (ISE) between the histogram and the model, a method rooted in the $L_{2}E$ theory proposed by scott2001parametric.
To align with the standard maximization notation of vandervaartwellner1996, we define the estimator $\widehat{\vartheta}_{n}$ as the maximizer of the negative squared distance. Let $\mathbb{M}_{n}(\vartheta)$ denote the empirical criterion and $\mathbb{M}(\vartheta)$ the population criterion:
@@
Here, $\widehat{f}_{n}(y)$ is the histogram density estimator and $f_{n}(y;\vartheta)$ is the discretized (binned) model density in (12). The population criterion $\mathbb{M}(\vartheta)$ corresponds to the negative squared $L_{2}$ distance between the true density $f_{0}$ and the candidate model density $f(\cdot;\vartheta)$. As discussed by scott2001parametric, maximizing $\mathbb{M}(\vartheta)$ is equivalent to minimizing the $L_{2}$ distance to the true data-generating distribution, which provides inherent robustness properties compared to maximum likelihood.
@@ 4.2 Model validation
5 Simulation study
@@
We evaluated performance across six sample sizes: $n\in\{100,500,1000,5000\}$. This range enables us to assess how the relative performance of $L_{2}$ and MLE changes with increasing information. For each combination of scenario and sample size, we generated 1,000 independent replicates. For each simulated dataset, we applied both the $L_{2}$ histogram-based estimator and MLE, recording parameter estimates and computation times for both methods. We construct histograms using Sturges’ rule to determine the number of bins: $k=\lceil\log_{2}(n)+1\rceil$. This binning rule yields between 8 bins ($n=100$) and 14 bins ($n=5000$).

### v3

@@ 1 Introduction
Serological data analysis plays a central role in reconstructing individual- and population-level exposure histories to infectious diseases, using information derived from antibody concentration measurements (Metcalf et al., 2016; Corran et al., 2007). Antibody levels can function as biomarkers of prior exposure, enabling the detection of both recent infections and long-term immune responses. In low-prevalence or post-elimination settings, serological data become particularly valuable as they can be used to identify subtle signs of disease resurgence or reintroduction (Drakeley et al., 2005; Arnold et al., 2018). Consequently, serological surveillance offers a powerful approach for monitoring disease dynamics, especially where traditional diagnostic methods may be insufficiently sensitive due to sporadic infections or predominantly asymptomatic transmission (Drakeley et al., 2005).
Gaussian mixture models (GMMs) are among the most widely used methods for analysing antibody concentration data (Fraley and Raftery, 2002; Arnold et al., 2018). In the serological context, a GMM typically assumes that the antibody concentration $y_{i}$ for the $i$-th individual arises from a mixture of two Gaussian distributions, representing seronegative and seropositive subpopulations:
@@
where $\pi_{0}+\pi_{1}=1$, and $(\mu_{k},\sigma_{k}^{2})$ denote the mean and variance of each component. This binary specification reflects the assumption that individuals fall into one of two mutually exclusive serological states, with no allowance for intermediate or uncertain exposure histories. Across a wide range of pathogens, such mixture models have become a standard way to operationalise immune activation from serological measurements by separating individuals into low- and high-response groups interpreted as immunologically naïve versus previously exposed or vaccinated. Although GMMs provide a flexible, data-driven approach to defining seropositivity thresholds, alternative strategies are also common in sero-epidemiology (Hay et al., 2024): using manufacturer-defined assay cutoffs; determining thresholds via ROC curves when positive and negative controls are available; basing cutoffs on the distribution of negative controls alone (typically two or three standard deviations above the mean); or, when available, applying immunological correlates of protection to define seroprotection-based thresholds.
@@
This approach builds on an extensive body of work applying GMMs to diverse pathogens for which serology provides a reliable marker of past exposure, including directly transmitted childhood infections such as measles, mumps and rubella (Vyse et al., 2006); parasitic and neglected tropical diseases such as onchocerciasis (Golden et al., 2016) and trachoma (Migchelsen et al., 2017); mosquito-borne infections such as malaria, where mixture models underpin estimates of age-specific seroprevalence (Sepúlveda et al., 2015); and viral infections characterised by substantial asymptomatic transmission such as hepatitis E (Katuwal et al., 2024) and SARS-CoV-2 (Bottomley et al., 2021). These applications highlight that the usefulness of serology strongly depends on the biology of the pathogen. As emphasised by Metcalf et al. (2016), serological surveys are most informative for infections in which antibody responses provide a durable and interpretable record of prior exposure or immunity. This includes infections that generate long-lasting protection, those with measurable antigenic variation, and those in which antibodies reliably reflect cumulative exposure even when infections are predominantly asymptomatic or clinically under-ascertained. The methods presented in this paper are relevant for the analysis of these types of pathogens.
@@
Recent methodological advances have sought to extend GMMs by incorporating covariates, such as age or other demographic factors, which influence antibody concentrations and seropositivity status. For example, Hardelid et al. (2008) used maternal age and country-of-birth to model variation in rubella antibody levels among newborns using a finite-mixture regression framework. Age-dependent mixture models are especially common in infectious disease applications: Del Fava et al. (2016) embedded a three-component mixture within a catalytic model to estimate age-specific immunity to varicella-zoster virus, while Cox et al. (2022) used penalised splines to let seropositivity probabilities vary smoothly with age in dengue serology data. Similarly, Kyomuhangi and Giorgi (2021) combined antibody acquisition and catalytic models to allow for temporally varying seroconversion and boosting rates. These approaches reflect cumulative exposure risk by allowing mixing proportions or transition probabilities to evolve with age or other covariates. In parallel, efforts have been made to relax the normality assumption within each component. A key extension replaces Gaussian components with skew-normal distributions (Azzalini, 1985), enabling the model to capture asymmetries commonly observed in antibody concentration data. Skew-normal mixtures retain analytical tractability while offering greater flexibility in representing the shape of seronegative and seropositive subpopulations (Dias Domingues et al., 2024).
@@
An alternative line of work has sought to avoid explicit dichotomisation through antibody acquisition (AA) models, in which age-dependent boosting and waning processes generate smooth trajectories of mean antibody levels (Yman et al., 2016). Although these models provide a useful mechanistic link between age, exposure, and antibody dynamics, they typically rely on a log-Gaussian assumption and do not adequately accommodate the skewness and multimodality that characterise many serological datasets. Semiparametric methods have also been proposed that use ensemble machine learning to estimate age-antibody curves nonparametrically (Arnold et al., 2017), but these target only the mean antibody level, precluding inference on other parameters of epidemiological interest. The proposed framework in this study unifies the strengths of these approaches by accommodating both mechanistic interpretability through AA-type formulations and flexible modeling of the full heterogeneous distribution of antibody levels across age through a latent variable formulation.
@@
A key practical challenge in implementing the proposed framework is the computational burden associated with maximum likelihood estimation, which requires repeated numerical integration. To address this, we develop a computationally efficient alternative based on histogram-based $L_{2}$ minimum distance estimation that substantially reduces the computational burden. Our work builds on the $L_{2}$ error theory developed by Scott (2001), which establishes the criteria under which $L_{2}$ estimators converge to the true parameter values. We extend this framework to our latent variable modelling framework and establish consistency of the resulting estimator under appropriate regularity conditions.
@@
The remainder of the paper is organised as follows. Section 2 introduces the proposed latent variable modelling framework, contrasts it with alternative latent formulations, and presents both single-density and mixture specifications for the latent seroreactivity distribution. Section 3 describes how age-dependent structure can be incorporated through mechanistic and data-driven parameterisations of both single Beta distributions and finite mixtures. Section 4 details maximum likelihood and histogram-based $L_{2}$ estimation procedures, establishes the consistency of the $L_{2}$ estimator, and describes model validation approaches. Section 5 presents two simulation studies: the first compares the statistical and computational efficiency of $L_{2}$ and maximum likelihood estimation, and the second evaluates the robustness of maximum likelihood estimation under model misspecification considering epidemiologically relevant data-generating mechanisms. Section 6 demonstrates the framework through applications to malaria serology data for AMA1 and MSP1 antigens, including age-stratified model comparisons. We conclude in Section 7 with methodological implications and extensions to multivariate settings, spatial-temporal modelling, and high-dimensional omics applications.
@@ 2 The proposed latent variable modelling framework for serological data analysis
To illustrate the interpretation, let us consider the simplest formulation of antibody acquisition models originally proposed by Yman et al. (2016). According to this model, we define
@@ 3.2 Finite mixture models with age dependent mixing probabilities
A wide range of mechanistic models for the relation between age and antibody levels has been developed in the context of infectious disease research. These include catalytic models, models that allow repeated boosting through multiple exposures, and models with time varying hazards; a comprehensive overview of these approaches can be found in Hay et al. (2024). Within our latent mixture framework, these formulations offer a convenient way to specify $\pi(a)$ while retaining epidemiological interpretability. For example, the catalytic framework assumes that the cumulative probability of activation follows
@@ 4 Inference
Parameter estimation for the latent variable model requires integrating over the continuous distribution of the unobserved immune state $T$, which presents computational challenges for likelihood-based inference. We address this by considering and comparing two approaches: standard maximum likelihood estimation (MLE), and a computationally efficient histogram-based $L_{2}$ distance minimization method. This can be thought of a discretizations of the $L_{2}E$ method proposed by Scott (2001). Although our main motivation for using the method is computational gain, the limiting case i.e the $L_{2}E$ method is also used for it’s robusntess Scott (2001).
@@
The theoretical justification for this approach rests on the fact that, as the bin widths $\Delta_{j}$ shrink, the histogram density $\widehat{f}_{j}$ converges to the empirical distribution of the data, while $f(m_{j};\vartheta,\psi)$ converges to the model density at those points. In this limiting regime, minimisation of (13) defines a minimum–distance estimator (Wolfowitz, 1957; Beran, 1977) that minimises the $L_{2}$ distance between the empirical and model densities, i.e. the $L_{2}E$ method.
@@ 4.1 Consistency of the histogram-based $L_{2}$ estimator
We establish the consistency of our estimator using the asymptotic framework of $M$-estimation (van der Vaart and Wellner, 1996). The estimator is defined by minimizing the integrated squared error (ISE) between the histogram and the model, a method rooted in the $L_{2}E$ theory proposed by Scott (2001).
To align with the standard maximization notation of van der Vaart and Wellner (1996), we define the estimator $\widehat{\vartheta}_{n}$ as the maximizer of the negative squared distance. Let $\mathbb{M}_{n}(\vartheta)$ denote the empirical criterion and $\mathbb{M}(\vartheta)$ the population criterion:
@@
Here, $\widehat{f}_{n}(y)$ is the histogram density estimator and $f_{n}(y;\vartheta)$ is the discretized (binned) model density in (12). The population criterion $\mathbb{M}(\vartheta)$ corresponds to the negative squared $L_{2}$ distance between the true density $f_{0}$ and the candidate model density $f(\cdot;\vartheta)$. As discussed by Scott (2001), maximizing $\mathbb{M}(\vartheta)$ is equivalent to minimizing the $L_{2}$ distance to the true data-generating distribution, which provides inherent robustness properties compared to maximum likelihood.
@@ 4.2 Model validation
5 Simulation studies
5.1 Assessment of the $L_{2}$ estimator performance
@@
We evaluated performance across four sample sizes: $n\in\{100,500,1000,5000\}$. This range enables us to assess how the relative performance of $L_{2}$ and MLE changes with increasing information. For each combination of scenario and sample size, we generated 1,000 independent replicates. For each simulated dataset, we applied both the $L_{2}$ histogram-based estimator and MLE, recording parameter estimates and computation times for both methods. We construct histograms using Sturges’ rule to determine the number of bins: $k=\lceil\log_{2}(n)+1\rceil$. This binning rule yields between 8 bins ($n=100$) and 14 bins ($n=5000$).
@@ 5.1.1 Results
5.2 Assessment of the impact of model misspesification
In this second simulation study, we assess the robustness of maximum likelihood estimation when the data-generating mechanism departs from the fitted model specification. Specifically, we consider three scenarios in which subpopulations experience different patterns of pathogen exposure that are not accounted for in the standard age-dependent structure.
The fitted model in all scenarios is the latent variable model with mean-variance parameterization of the latent Beta distribution as given in equation (8), where the mean $\mu(a)$ depends on age through the logit-linear specification in (9). The true parameters governing the baseline age-dependent distribution are $\eta_{0}=-2.0$, $\eta_{1}=0.4$, and $\phi=3.0$ for the distribution of $T$, and $\mu_{0}=-3.5$, $\mu_{1}=1.0$, $\sigma_{0}=0.8$, and $\sigma_{1}=0.3$ for the conditional distribution of $Y$ given $T$. Rather than sampling directly from this model, we introduce structured heterogeneity representing plausible departures from the assumed age-dependent dynamics. For each scenario, we consider sample sizes $n\in\{1000,2000,5000\}$ and generate 2000 simulated data-sets.
The first scenario represents populations in which a proportion of individuals have received interventions that directly elevate antibody levels through mechanisms distinct from natural infection. Examples include prophylactic antibody therapy or vaccines that produce measurable serological responses that would be reflected in elevated values of the latent ser-reactivity state. Hence, we simulate under this scenario by first sampling each individual’s baseline latent state $T_{i}$ from the age-dependent Beta distribution, then adding a stochastic boost $\Delta_{i}\sim\text{Beta}(3,3)$ to a randomly selected subset representing vaccinated individuals. The boosted latent state is given by $T_{i}^{*}=\min(1,T_{i}+\Delta_{i})$, where the truncation at unity reflects the upper bound of the latent scale. We shall use the parameter $\rho\in\{0.2,0.5,0.8\}$ to denote the coverage of the intervention, assumed to be uniform across all age groups.
The second scenario represents age-targeted public health interventions that indirectly reduce pathogen exposure for specific age groups. Examples include school-based health programs, nutritional supplementation initiatives, or targeted vector control efforts that primarily benefit children of school age. Rather than directly boosting antibody levels, these interventions reduce the force of infection, thereby dampening the age-dependent accumulation of seroreactivity. We implement this by modifying the expected latent seroreactivity for individuals aged between 5 and 15 years according to
$\mu^{*}(a)=\frac{\mu(a)\exp(-\delta)}{\mu(a)\exp(-\delta)+1-\mu(a)},$
where $\mu(a)$ denotes the age-specific mean under the standard model and $\delta\in\{0.4,0.7,1.2\}$ quantifies the strength of exposure reduction. Larger values of $\delta$ correspond to greater reductions in the expected seroreactivity for the targeted age range. For ages outside the interval $[5,15]$, we set $\mu^{*}(a)=\mu(a)$. The latent state for each individual is then sampled from $T_{i}\sim\text{Beta}(\mu^{*}(a_{i})\phi,[1-\mu^{*}(a_{i})]\phi)$.
The third scenario represents immigration or population movement that introduces individuals with no prior pathogen exposure into endemic areas. In cross-sectional serological surveys, such individuals may be sampled before they have had sufficient contact with the pathogen to mount a detectable immune response, resulting in an absence of measurable seroreactivity at the time of sampling. We model this by designating a proportion $r\in\{0.10,0.25,0.40\}$ of individuals as completely unexposed, assigning them $T_{i}=0$ exactly, while the remaining 1r1-r proportion are sampled from the age-dependent Beta distribution. This creates a discrete point mass at zero seroreactivity superimposed on the continuous age-dependent distribution, representing a fundamentally different form of population heterogeneity than the model assumes.
After fitting the mean-variance Beta model to each of the 2,000 simulated datasets, we compute the median bias and the median absolute deviation for parameters of the conditional distribution of $Y_{i}$, namely $\mu_{0}$, $\mu_{1}$, $\sigma^{2}_{0}$ and $\sigma^{2}_{1}$, and the parameters of the latent variable $T_{i}$, consisting of $\eta_{0}$, $\eta_{1}$ and the variance parameter $\phi$.
Although we monitor estimation performance across all seven model parameters, our primary interest lies in the recovery of $\eta_{0}$ and $\eta_{1}$, which govern the age-dependent mean of the latent seroreactivity distribution and therefore determine the epidemiological inferences drawn from the fitted model.
5.2.1 Results
Table 4: Scenario 1: Intervention-induced antibody elevation. Parameter $\rho$ represents intervention coverage. Median bias and median absolute error (MAE) are reported for all seven model parameters across sample sizes and coverage levels. Results based on 2000 simulation replicates per configuration.
Parameter $n$ $\rho=0.2$ $\rho=0.5$ $\rho=0.8$
Bias MAE Bias MAE Bias MAE
$\mu_{0}$ 1000 0.416 0.542 0.510 0.589 0.618 0.670
2000 0.276 0.367 0.383 0.426 0.545 0.558
5000 0.227 0.271 0.352 0.361 0.522 0.523
$\mu_{1}$ 1000 -0.105 0.569 0.136 0.467 0.220 0.387
2000 0.129 0.386 0.289 0.386 0.299 0.349
5000 0.175 0.260 0.317 0.334 0.327 0.334
$\sigma_{0}$ 1000 0.065 0.113 0.079 0.119 0.084 0.119
2000 0.045 0.079 0.063 0.084 0.076 0.091
5000 0.039 0.053 0.059 0.065 0.075 0.076
$\sigma_{1}$ 1000 -0.015 0.300 -0.133 0.300 -0.192 0.300
2000 -0.078 0.300 -0.176 0.300 -0.184 0.255
5000 -0.068 0.133 -0.145 0.161 -0.164 0.170
$\eta_{0}$ 1000 -0.041 0.594 0.021 0.489 0.079 0.417
2000 0.099 0.326 0.138 0.276 0.143 0.248
5000 0.152 0.194 0.162 0.182 0.166 0.183
$\eta_{1}$ 1000 -0.104 0.135 -0.096 0.114 -0.077 0.097
2000 -0.084 0.089 -0.078 0.082 -0.067 0.071
5000 -0.070 0.070 -0.069 0.069 -0.064 0.064
$\phi$ 1000 -0.683 1.352 -0.595 0.994 -0.607 0.839
2000 -0.227 0.693 -0.234 0.531 -0.429 0.547
5000 -0.083 0.386 -0.137 0.300 -0.335 0.376
Under Scenario 1 (Table 4), the most notable consequence of the intervention-induced boost is a systematic upward displacement of $\mu_{0}$, the lower boundary parameter of the conditional distribution. The median bias in $\mu_{0}$ increases monotonically with coverage $\rho$, ranging from $0.42$ at $\rho=0.2$ to $0.62$ at $\rho=0.8$ for $n=1000$, and remaining non-negligible even at $n=5000$. Importantly, however, the age-trend parameters $\eta_{0}$ and $\eta_{1}$ show small median bias across all sample sizes and $\rho$ values, with median biases not exceeding $0.17$ and $0.10$ in absolute value respectively at $n=5000$. The precision parameter $\phi$ shows moderate negative bias at small sample sizes that largely decreases by $n=5000$. The parameter $\sigma_{1}$ exhibits MAE of around $0.3$ at $n=1000$ and $n=2000$, indicating that the high-seroreactivity variance becomes poorly identified when the upper tail of the antibody distribution is distorted by the boost, though this also decreases at $n=5000$. Overall, the observed patterns can be explained by the fact that the random boosting process shifts the observed antibody distribution upward, and the model largely absorbs part of this shift into the lower boundary estimate of the conditional model of $Y_{i}$ rather than into the age-dependent latent structure of $T_{i}$.
Table 5: Scenario 2: Age-targeted exposure reduction for individuals aged 5–15 years. Parameter $\delta$ represents the strength of exposure dampening on the logit scale for the age group 5–15 years. Median bias and MAE reported for all seven model parameters. Results based on 2000 simulation replicates per configuration.
Parameter $n$ $\delta=0.4$ $\delta=0.7$ $\delta=1.2$
Bias MAE Bias MAE Bias MAE
$\mu_{0}$ 1000 0.410 0.456 0.371 0.408 0.299 0.323
2000 0.356 0.375 0.334 0.347 0.268 0.276
5000 0.312 0.318 0.307 0.312 0.255 0.258
$\mu_{1}$ 1000 -0.688 0.890 -0.755 0.968 -0.626 0.902
2000 -0.293 0.375 -0.283 0.354 -0.282 0.348
5000 -0.205 0.242 -0.218 0.244 -0.193 0.230
$\sigma_{0}$ 1000 0.075 0.095 0.074 0.090 0.072 0.085
2000 0.071 0.081 0.077 0.083 0.074 0.079
5000 0.066 0.068 0.075 0.077 0.073 0.075
$\sigma_{1}$ 1000 0.181 0.300 0.214 0.300 0.179 0.300
2000 0.061 0.300 0.062 0.300 0.064 0.300
5000 0.042 0.114 0.047 0.123 0.043 0.116
$\eta_{0}$ 1000 -0.972 1.116 -1.463 1.543 -1.804 1.827
2000 -0.612 0.643 -1.003 1.017 -1.451 1.451
5000 -0.469 0.474 -0.866 0.871 -1.324 1.324
$\eta_{1}$ 1000 0.023 0.292 0.134 0.442 0.270 0.517
2000 0.070 0.154 0.177 0.256 0.311 0.391
5000 0.092 0.110 0.201 0.219 0.337 0.355
$\phi$ 1000 -1.316 2.311 -1.380 2.566 -1.151 2.540
2000 -0.749 0.842 -0.733 0.806 -0.671 0.734
5000 -0.588 0.605 -0.619 0.640 -0.548 0.570
Scenario 2 produces the most severe and structurally persistent misspecification of the three scenarios. The suppression of seroreactivity among individuals aged 5 to 15 years creates a local dip in the age-antibody profile that the fitted logit-linear model in age cannot represent, resulting in a systematic distortion of the age-trend parameters. Specifically, $\eta_{0}$ shows large negative median bias that worsens monotonically with $\delta$ and fails to diminish meaningfully with increasing sample size: at $n=5000$, the median bias reaches 0.47-0.47, 0.87-0.87, and 1.32-1.32 for $\delta=0.4$, $0.7$, and $1.2$ respectively, corresponding to relative errors of up to $66\%$ on a parameter whose true value is 2.0-2.0. The parameter $\eta_{1}$ acquires positive median bias that also increases with $\delta$ and persists at $n=5000$ ($0.09$, $0.20$, $0.34$), indicating that the model compensates for the artificially low intercept by steepening the estimated age gradient. This compensation mechanism is internally coherent but epidemiologically misleading: the fitted model would suggest a steeper accumulation of seroreactivity with age than is actually present in the unaffected age groups. The lower boundary $\mu_{0}$ also acquires persistent positive bias of approximately $0.3$ that does not diminish with $n$, while $\phi$ shows sustained negative bias reflecting reduced apparent precision when the latent age structure is misspecified. The non-vanishing nature of the biases in $\eta_{0}$ and $\eta_{1}$ as $n$ increases confirms that this scenario induces genuine structural misspecification that cannot be resolved through larger sample sizes alone.
Table 6: Scenario 3: Introduction of immunologically naive subpopulation. Parameter $r$ represents the proportion of unexposed individuals. Median bias and MAE reported for all seven model parameters. Results based on 2000 simulation replicates per configuration.
Parameter $n$ $r=0.10$ $r=0.25$ $r=0.40$
Bias MAE Bias MAE Bias MAE
$\mu_{0}$ 1000 0.309 0.392 0.183 0.246 0.091 0.156
2000 0.232 0.293 0.154 0.190 0.066 0.120
5000 0.188 0.202 0.125 0.135 0.056 0.083
$\mu_{1}$ 1000 -0.635 0.752 -1.013 1.068 -1.229 1.270
2000 -0.331 0.415 -0.653 0.675 -0.800 0.802
5000 -0.282 0.314 -0.587 0.589 -0.732 0.732
$\sigma_{0}$ 1000 0.075 0.101 0.057 0.079 0.036 0.059
2000 0.067 0.081 0.061 0.068 0.036 0.045
5000 0.062 0.064 0.056 0.057 0.036 0.038
$\sigma_{1}$ 1000 0.177 0.300 0.296 0.300 0.369 0.369
2000 0.081 0.300 0.203 0.300 0.246 0.300
5000 0.076 0.137 0.189 0.237 0.243 0.294
$\eta_{0}$ 1000 -0.118 0.583 -0.086 0.555 -0.190 0.513
2000 0.072 0.308 0.082 0.328 0.009 0.296
5000 0.133 0.190 0.185 0.226 0.115 0.192
$\eta_{1}$ 1000 -0.115 0.156 -0.157 0.190 -0.192 0.225
2000 -0.097 0.104 -0.129 0.137 -0.161 0.167
5000 -0.077 0.077 -0.099 0.099 -0.124 0.126
$\phi$ 1000 -1.270 1.502 -1.741 1.866 -1.994 2.053
2000 -0.885 1.006 -1.415 1.435 -1.636 1.651
5000 -0.764 0.803 -1.315 1.316 -1.567 1.567
In Scenario 3, the primary consequence of naive immigration is a systematic underestimation of the saturation parameter $\mu_{1}$, the upper boundary of the conditional distribution. The median bias in $\mu_{1}$ worsens progressively with the proportion of unexposed individuals $\pi$, reaching 0.64-0.64, 1.01-1.01, and 1.23-1.23 at $n=1000$ for $\pi=0.10$, $0.25$, and $0.40$ respectively. Although these biases decrease with ample size, they remain substantial at $n=5000$ (0.28-0.28, 0.59-0.59, 0.73-0.73), reflecting the fundamental incompatibility between the point mass at $T_{i}=0$ for naive individuals and the continuous Beta distribution assumed by the model. A closely related pattern appears in $\phi$, which shows large negative median bias that also worsens with $r$ and persists at large $n$ (0.76-0.76, 1.32-1.32, 1.57-1.57 at $n=5000$). This is because the excess mass near $T_{i}=0$ makes the latent distribution appear more diffuse than the fitted Beta can represent, so the precision parameter is consistently underestimated (we recall that lower values in $\phi$ increse the variance of $T$). By contrast, the age-trend parameters $\eta_{0}$ and $\eta_{1}$ exhibit only modest median biases that decrease toward zero as $n$ increases, suggesting that the age-dependent structure is approximately recoverable even in the presence of naive immigration, provided the sample is large enough.
Comparing across the three scenarios, a clear hierarchy of severity emerges. Scenario 2 is the most damaging because it induces structural misspecification in the age-trend parameters $\eta_{0}$ and $\eta_{1}$ that persists regardless of sample size, rendering the epidemiological interpretation of the fitted model unreliable even when $n=5000$. Scenario 3 occupies an intermediate position: while the boundary and precision parameters are substantially distorted, the age-trend parameters remain approximately recoverable at large sample sizes, and the direction of the bias in $\mu_{1}$ provides a diagnostically useful signal of unexposed subpopulation presence. Scenario 1 is the mildest departure, with misspecification confined primarily to the lower boundary parameter $\mu_{0}$ while leaving the age-dependent structure intact. These findings suggest that the proposed modelling framework is most sensitive to misspecification that disrupts the global age-trend structure, and most robust to departures that act uniformly across age, which are largely absorbed into the estimated parameters of the conditional distribution of $Y_{i}$ given $T_{i}$.